Haemoglobin contains 0.33% of Iron by weight. The molecular weight of it is approx. 67200. The no. of iron atoms present in one molecule of Haemoglobin are ________?


Options :


(1) 1 (2) 2 (3) 4 (4) 6

Grade:11

5 Answers

Vikas TU
10 years ago

Ist method:

a BE THE NO. OF ATOMS PRESENT IN ONE MOLECULE:

(a/67200) x 100 = 0.33 %

a = 3.9 = 4 (approx)

 

 

2nd method (direct method)

1) 100 - 0.33 grams of haemoglobin is present then,

moles of haemoglobin = 99.67/67200 = 0.00148

 

2) Moles of iron present in haemo. = 0.33/56 = 0.00589

 

Now, a/67200 moles of Haemoglobin contains  0.00589  moles
 then  no. of moles of iron present in one molecule of haemo. =  0.00589 x Na/0.00148 x Na = 3.9 = 4

 

Approve plz!

Amruth Sai
10 years ago

(3)4

Abhishek kumar singh
6 years ago
Weight of fe in heamoglobin = (0.33 ×67200) ÷100=672÷3=224Moles of fe = wt. In gm÷molar mass=224÷56=4
Anurag Singh
5 years ago
Anurag singh 
Hello friends ,
The hemoglobin molecular mass=67200.
0.33%of iron = given mass ×100/molar mass
0.33% of iron =given mass×100/67200
Given mass =221
Iron mass =221/56
Iron atoms =4 
Thank u 
 
ankit singh
3 years ago
The number of iron atoms (At. Wt. of Fe=56) present in one molecule of haemoglobin is. Gram atom of Fe=672×0.3356=4.